# Scaling Inference-Efficient Language Models

### v1

@@ Abstract
Scaling laws are powerful tools to predict the performance of large language models. However, current scaling laws fall short of accounting for inference costs. In this work, we first show that model architecture affects inference latency, where models of the same size can have up to 3.5×3.5$\times$3.5 × difference in latency. To tackle this challenge, we modify the Chinchilla scaling laws to co-optimize the model parameter count, the number of training tokens, and the model architecture. Due to the reason that models of similar training loss exhibit gaps in downstream evaluation, we also propose a novel method to train inference-efficient models based on the revised scaling laws. We perform extensive empirical studies to fit and evaluate our inference-aware scaling laws. We vary model parameters from 80M to 1B, training tokens from 1.6B to 30B, and model shapes, training a total of 63 models. Guided by our inference-efficient scaling law and model selection method, we release the Morph-1B model, which improves inference latency by 1.8×1.8$\times$1.8 × while maintaining accuracy on downstream tasks compared to open-source models, pushing the Pareto frontier of accuracy-latency tradeoff.
@@ 1 Introduction
In this work, we first show that the number of parameters is not the exclusive factor affecting inference efficiency. As illustrated in Figure 2, the model architecture also plays a critical role. Following this observation, we introduce inference-efficient scaling laws, building upon the Chinchilla scaling law and incorporating model architecture considerations. Additionally, due to the disparity between model loss and accuracy in downstream tasks, we develop a novel method (Figure 6) that utilizes inference-efficient scaling laws to rank various model architectural choices. Our findings suggest that the relative ranking of loss predictions from scaling laws is more significant than their absolute values (§2).
@@
Lastly, we train the Morph-1B model using the best model configuration predicted by our inference-efficient scaling law and ranking algorithm. Figure 1 summarizes our main results. Compared to other open source models of similar size, Morph-1B improves the inference latency by 1.8×1.8$\times$1.8 × while maintaining accuracy over downstream tasks. These findings underscore the effectiveness of our inference-efficient scaling law. By designing a general scaling law that focuses on inference latency, our work can also capture the accuracy-efficiency trade-off for recent and future architectural optimizations, such as GQA (Touvron et al., 2023b; Dubey et al., 2024) and MLA (Liu et al., 2024a).
@@
Figure 3: Model Shape on Inference Latency: (Left) We illustrate the correlation between inference latency and the number of layers, with constant hidden size. Due to the sequential nature of LLM execution, latency increases linearly with the number of layers. (Center) We plot the relationship between inference latency and hidden size with the number of layers fixed. We see that model width does not affect latency for smaller models but only for larger models. (Right) We show the relationship between inference latency and aspect ratio, with the number of model parameters fixed. We see a downward trend in inference latency as we make the model wider and shallower.
@@ 2.2 Inference-Efficient Scaling Laws
Motivated by Figure 2, we closely examine the effect of model configuration on inference latency by altering the hidden size $d_{\text{model}}$ and number of layers $n_{\text{layers}}$ as shown in Figure 3.
Figure 3(a) shows that inference latency increases linearly with the number of layers when the hidden size remains constant. This occurs as the inference computation must be performed sequentially, one layer at a time (Yan et al., 2024). However, the matrix computations within a single layer can be performed in parallel. Furthermore, Figure 3(c) indicates that for the same number of parameters (say 7B), we can achieve different latency targets by changing the ratio of the number of hidden parameters in one layer ($d_{\text{model}}$) vs. the number of layers ($n_{\text{layers}}$).
@@ 3.1 Experimental Setup
Furthermore, to compare the predicted loss against the actual loss, we measure relative prediction error: $|\psi-\hat{\psi}|/\psi$, mean squared error (MSE): $\frac{1}{n}\sum_{i=1}^{n}(\psi_{i}-\hat{\psi}_{i})$, and $R^{2}=1-\sum_{i=1}^{n}(\psi_{i}-\hat{\psi}_{i})^{2}/\sum_{i=1}^{n}(\psi_{i}-\bar{\psi})^{2}$, where $\psi$ represents the actual loss, $\hat{\psi}$ the predicted loss from scaling laws, and $\bar{\psi}=\frac{1}{n}\sum_{i=1}^{n}\psi_{i}$. We also apply Spearman’s rank correlation coefficient (Spearman, 1961) to evaluate how well the predicted rankings correspond to the actual rankings.
@@ 4 Results
Table 1: Data Used to Fit Scaling Laws: In this table, we show the number of parameters and tokens used in model training to fit the scaling laws in Figure 7-9. ✓ indicates we use all model variants with the given size and ✗ means we do not use any model variants with the given size. ❖ indicates that we randomly sample one model variant from the candidate set. The details of model variants are included in Appendix A.
@@ 4.1 Prediction acccuracy
Table 2: Inference-Efficient Models: In this table, we compare the results of of Morph-1B variants against other open pretrained models of similar size. The evaluation of large language models such as Open-LM-1B (Gururangan et al., 2023), OPT-1.3B (Zhang et al., 2022), Pythia-1.3B (Biderman et al., 2023), Neox-1.3B (Black et al., 2022) and OPT-IML-1.3B (Iyer et al., 2022) is summarized from (Gururangan et al., 2023).
@@ 5 Related Work
Scaling Laws. Scaling laws are powerful predictors for how large language models behave as parameters increase (Kaplan et al., 2020). Plenty of subsequent works have contributed to the development of scaling laws (Hoffmann et al., 2022; Muennighoff et al., 2023; Sardana et al., 2023; Tao et al., 2024; Kumar et al., 2024; Gadre et al., 2024; Ruan et al., 2024; Abnar et al., 2025). In particular, Chinchilla scaling law (Hoffmann et al., 2022) optimizes a fixed computing budget allocation by balancing the number of model parameters against the number of training tokens to minimize the training loss. Data-Constrained scaling law (Muennighoff et al., 2023) extends the Chinchilla scaling laws by considering repeated data. The scaling laws presented in (Gadre et al., 2024) not only predict training loss under over-training scenarios but also connect training loss to downstream error. Beyond Chinchilla-Optimal (Sardana et al., 2023) attempted to account for inference cost in their scaling law. However, unlike training tokens, the number of inference tokens cannot be measured in advance. Our paper proposes a scaling law that avoids estimating the tokens generated during inference.
Inference Serving. Inference cost has drawn significant attention in recent years. Many inference systems and algorithms have been developed to speed up model serving (Olston et al., 2017; Gujarati et al., 2020; Gugger et al., 2022; Yu et al., 2022; Leviathan et al., 2023; Kwon et al., 2023; Zheng et al., 2023; Agarwal et al., 2024a, b; Ye et al., 2025; MLC team, 2023-2025). Specifically, Orca (Yu et al., 2022) utilizes continuous batching to achieve higher inference throughput. vLLM (Kwon et al., 2023) improves the throughput of popular LLMs by using PagedAttention to manage the KV cache memory. Furthermore, SGLang (Zheng et al., 2023) improves the inference throughput and latency by using RadixAttention. A recent study introduces FlashInfer (Ye et al., 2025), which employs block-sparse and composable formats to tackle KV cache storage heterogeneity.
@@ Appendix B Results over A30 GPUs
Appendix C Results over vLLM
@@
Appendix D More Scaling Laws Fits
@@
Appendix E Evaluation Dataset Details
@@
Appendix F Contribution Statement

### v2

@@ Abstract
Scaling laws are powerful tools to predict the performance of large language models. However, current scaling laws fall short of accounting for inference costs. In this work, we first show that model architecture affects inference latency, where models of the same size can have up to 3.5×3.5$\times$3.5 × difference in latency. To tackle this challenge, we modify the Chinchilla scaling laws to co-optimize the model parameter count, the number of training tokens, and the model architecture. Due to the reason that models of similar training loss exhibit gaps in downstream evaluation, we also propose a novel method to train inference-efficient models based on the revised scaling laws. We perform extensive empirical studies to fit and evaluate our inference-aware scaling laws. We vary model parameters from 80M to 1B, training tokens from 1.6B to 30B, and model shapes, training 63 models. Guided by our inference-efficient scaling law and model selection method, we release the Morph-1B model, which improves inference latency by 1.8×1.8$\times$1.8 × while maintaining accuracy on downstream tasks compared to open-source models, pushing the Pareto frontier of accuracy-latency tradeoff. Notably, our experiments reveal that wider and shallower models can yield efficiency gains while preserving accuracy.
@@ 1 Introduction
We first show that the number of parameters is not the exclusive factor affecting inference efficiency. As illustrated in Figure 2, the model architecture also plays a critical role. Following this observation, we introduce inference-efficient scaling laws, building upon the Chinchilla scaling law and incorporating model architecture considerations. Additionally, due to the disparity between model loss and accuracy in downstream tasks, we develop a novel method (Figure 6) that utilizes inference-efficient scaling laws to rank various model architectural choices. Our findings suggest that the relative ranking of loss predictions from scaling laws is more significant than their absolute values (§2).
@@
Lastly, we train the Morph-1B [^2] The training code is available at https://github.com/Waterpine/open-lm-morph. The Morph-1B model checkpoint is available at https://huggingface.co/NaiveUser/morph-1b. model using the best model configuration predicted by our inference-efficient scaling law and ranking algorithm. Figure 1 summarizes our main results. Compared to other open source models of similar size, Morph-1B improves the inference latency by 1.8×1.8$\times$1.8 × while maintaining accuracy over downstream tasks. These findings underscore the effectiveness of our inference-efficient scaling law. By designing a general scaling law that focuses on inference latency, our work can also capture the accuracy-efficiency trade-off for recent and future architectural optimizations, such as GQA (Touvron et al., 2023b; Dubey et al., 2024) and MLA (Liu et al., 2024a).
@@
Figure 3: Model Shape on End-to-End Inference Latency: (Left) We illustrate the correlation between inference latency and the number of layers, with constant hidden size. Due to the sequential nature of LLM execution, latency increases linearly with the number of layers. (Center) We plot the relationship between inference latency and hidden size with the number of layers fixed. We see that model width does not affect latency for smaller models but only for larger models. (Right) We show the relationship between inference latency and aspect ratio, with the number of model parameters fixed. We see a downward trend in inference latency as we make the model wider and shallower. All evaluations were performed using the Hugging Face generate function on a single NVIDIA Ampere 40GB A100 GPU with batch size 1, input length 128, and output length 256.
@@ 2.1 Preliminaries
(a) 1B Model Variants
(b) 3B Model Variants
(c) 7B Model Variants
Figure 4: Model Shape on Throughput: We examine the relationship between inference throughput and model architecture by fixing the total parameter count and varying the hidden size and number of layers. Across different batch sizes, wider and shallower models consistently yield better inference throughput for large language models. Each tuple in the legend represents a model configuration: the first number is the hidden size $d_{\text{model}}$, and the second is the number of layers $n_{\text{layers}}$. All evaluations were performed using the Hugging Face generate function on a single NVIDIA Ampere 40GB A100 GPU with input length 128, and output length 256.
@@ 2.2 Inference-Efficient Scaling Laws
Motivated by Figure 2, we closely examine the effect of the aspect ratio ($d_{\text{model}}/n_{\text{layers}}$) on inference latency and throughput by altering the hidden size $d_{\text{model}}$ and the number of layers $n_{\text{layers}}$ as shown in Figure 3 and Figure 4. Reasonable aspect ratios are chosen based on open-weight models listed in Appendix G.
(a) PIQA
(b) BoolQ
(c) HellaSwag
Figure 3(a) shows that inference latency increases linearly with the number of layers when the hidden size remains constant. This occurs as the inference computation must be performed sequentially, one layer at a time (Yan et al., 2024). However, the matrix computations within a single layer can be performed in parallel. Furthermore, Figure 3(c) indicates that for the same number of parameters, we can achieve different latency targets by changing the ratio of the number of hidden parameters in one layer ($d_{\text{model}}$) vs. the number of layers ($n_{\text{layers}}$). Moreover, in Figure 4, We study the relationship between model shape and inference throughput under a fixed parameter budget. We observe that, under a fixed parameter budget, wider and shallower models consistently achieve higher inference throughput. Due to space constraints, results on the relationship between aspect ratio and time to first token (TTFT) are provided in Appendix C.
@@ 3.1 Experimental Setup
Furthermore, to compare the predicted loss against the actual loss, we measure relative prediction error: $|\psi-\hat{\psi}|/\psi$, mean squared error (MSE): $\frac{1}{n}\sum_{i=1}^{n}(\psi_{i}-\hat{\psi}_{i})^{2}$, and $R^{2}=1-\sum_{i=1}^{n}(\psi_{i}-\hat{\psi}_{i})^{2}/\sum_{i=1}^{n}(\psi_{i}-\bar{\psi})^{2}$, where $\psi$ represents the actual loss, $\hat{\psi}$ the predicted loss from scaling laws, and $\bar{\psi}=\frac{1}{n}\sum_{i=1}^{n}\psi_{i}$. We also apply Spearman’s rank correlation coefficient (Spearman, 1961) to evaluate how well the predicted rankings correspond to the actual rankings.
@@ 4 Results
Table 1: Data Used to Fit Scaling Laws: In this table, we show the number of parameters and tokens used in model training to fit the scaling laws in Figure 8-10. ✓ indicates we use all model variants with the given size and ✗ means we do not use any model variants with the given size. ❖ indicates that we randomly sample one model variant from the candidate set. The details of model variants are included in Appendix A.
@@ 4.1 Prediction acccuracy
Table 2: Inference-Efficient Models: In this table, we compare the results of Morph-1B variants against other open pretrained models of similar size. The evaluation of large language models such as Open-LM-1B (Gururangan et al., 2023), OPT-1.3B (Zhang et al., 2022), Pythia-1.3B (Biderman et al., 2023), Neox-1.3B (Black et al., 2022) and OPT-IML-1.3B (Iyer et al., 2022) is summarized from (Gururangan et al., 2023).
@@ 5 Related Work
Scaling Laws. Scaling laws are powerful predictors for how large language models behave as parameters increase (Kaplan et al., 2020). Plenty of subsequent works have contributed to the development of scaling laws (Hoffmann et al., 2022; Muennighoff et al., 2023; Sardana et al., 2023; Tao et al., 2024; Kumar et al., 2024; Gadre et al., 2024; Ruan et al., 2024; Abnar et al., 2025; Krajewski et al., 2024). In particular, Chinchilla scaling law (Hoffmann et al., 2022) optimizes a fixed computing budget allocation by balancing the number of model parameters against the number of training tokens to minimize the training loss. Data-Constrained scaling law (Muennighoff et al., 2023) extends the Chinchilla scaling laws by considering repeated data. The scaling laws presented in (Gadre et al., 2024) not only predict training loss under over-training scenarios but also connect training loss to downstream error. Beyond Chinchilla-Optimal (Sardana et al., 2023) attempted to account for inference cost in their scaling law. However, unlike training tokens, the number of inference tokens cannot be measured in advance.
Inference Serving Systems. Inference cost has drawn significant attention in recent years. Many inference systems and algorithms have been developed to speed up model serving (Olston et al., 2017; Gujarati et al., 2020; Gugger et al., 2022; Yu et al., 2022; Leviathan et al., 2023; Kwon et al., 2023; Zheng et al., 2023; Agarwal et al., 2024a, b; Ye et al., 2025; MLC team, 2023-2025). Specifically, Orca (Yu et al., 2022) utilizes continuous batching to achieve higher inference throughput. vLLM (Kwon et al., 2023) improves the throughput of popular LLMs by using PagedAttention to manage the KV cache memory. Furthermore, SGLang (Zheng et al., 2023) improves the inference throughput and latency by using RadixAttention. A recent study introduces FlashInfer (Ye et al., 2025), which employs block-sparse and composable formats to tackle KV cache storage heterogeneity.
Compute-Efficient Model Design. Previous research has explored the trade-offs of various model configurations in Vision Transformers (ViTs) (Alabdulmohsin et al., 2023). Additionally, (Tay et al., 2021) demonstrates that training deep and narrow models can be particularly beneficial when computational resources are limited. More recently, several efficient attention mechanisms (Xiao et al., 2023; Gao et al., 2024; Jiang et al., 2024; Xiao et al., 2024; Yuan et al., 2025) have been introduced to enhance inference efficiency by modifying the attention block.
@@ Appendix B Results over A30 GPUs
Appendix C More Results over A100 GPUs
In this section, we evaluate the relationship between model architecture and Time To First Token (TTFT) over a single NVIDIA Ampere 40GB A100 GPU by fixing the total parameter count and varying the hidden size and number of layers. From Figure 14, we observe that wider and shallower models consistently achieve lower TTFT.
(a) 1B Model Variants
(b) 3B Model Variants
(c) 7B Model Variants
Figure 14: Model Shape on Time To First Token (TTFT): We examine the relationship between TTFT and model architecture by fixing the total parameter count and varying the hidden size and number of layers. Across different batch sizes, wider and shallower models consistently achieve lower TTFT. Each tuple in the legend represents a model configuration: the first number is the hidden size $d_{\text{model}}$, and the second is the number of layers $n_{\text{layers}}$. All evaluations were performed using the Hugging Face generate function on a single NVIDIA Ampere 40GB A100 GPU with input length 128, and output length 1.
Appendix D Results over vLLM
@@
Appendix E More Scaling Laws Fits
@@
Appendix F Evaluation Dataset Details
@@
Appendix G Open-Source Model Configurations
In this section, Table 6 presents model configurations from Hugging Face, highlighting the vast space of architectural design choices.
Table 6: Model Configurations: We present the configurations of models available on Hugging Face.
Model $d_{\text{model}}$ $n_{\text{layers}}$ $d_{\text{model}}$ / $n_{\text{layers}}$
Llama-3.2-1B (Dubey et al., 2024) 2048 16 128
Llama-3.2-3B (Dubey et al., 2024) 3072 28 109.7
Qwen2.5-0.5B (Yang et al., 2024) 896 24 37.3
Qwen2.5-1.5B (Yang et al., 2024) 1536 28 54.9
Qwen2.5-3B (Yang et al., 2024) 2048 36 56.9
Qwen2.5-7B (Yang et al., 2024) 3584 28 128
Qwen2.5-14B (Yang et al., 2024) 5120 48 106.7
gemma-2b (Team et al., 2024a) 2048 18 113.8
gemma-7b (Team et al., 2024a) 3072 28 109.7
gemma-2-2b (Team et al., 2024b) 2304 26 88.6
gemma-2-9b (Team et al., 2024b) 3584 42 85.3
gemma-2-27b (Team et al., 2024b) 4608 46 100.2
microsoft-phi-2 (Phi, 2023) 2560 32 80
microsoft-phi-4 (Abdin et al., 2024) 5120 40 128
Appendix H Parameter Fits
Table 7: Parameter Fits: Coefficients for the scaling laws presented in Figure 8.
Law A B E $\alpha$ $\epsilon$
Chinchilla 7720.62 68572.73 2.13 0.49 /
Inference-Efficient 54754.14 778340.38 2.45 0.61 0.0011
Table 8: Parameter Fits: Coefficients for the scaling laws presented in Figure 9.
Law A B E $\alpha$ $\epsilon$
Chinchilla -25287.67 248461.43 2.14 0.51 /
Inference-Efficient -16247.15 958437.97 2.41 0.60 0.0011
Table 9: Parameter Fits: Coefficients for the scaling laws presented in Figure 10.
Law A B E $\alpha$ $\epsilon$
Chinchilla 793.45 4090.85 1.09 0.35 /
Inference-Efficient 32515.16 408925.99 2.34 0.58 0.0016
Appendix I Contribution Statement